OpenStudy (anonymous):

I am just wondering whether there is an more efficient way than trial division on finding out inverse of mod ( that is , given positive integer a, p, find an positive b, 0<= b <= p which leads a * b = 1 (mod p)) ? sorry for my poor English :(

OpenStudy (anonymous):

if a and p are relatively prime then \[a^{p-1}\equiv 1 (\text{mod} p)\]

OpenStudy (anonymous):

so to solve \[ab\equiv 1 (\text{mod} p)\] for b you can take \[b=a^{p-2}\] i guess. with small numbers trial and error may work best.

OpenStudy (anonymous):

for example if i want to solve \[6b\equiv 1 (\text{mod} 11)\] i could take \[6^{9}\] lets check \[6^2=3, 6^3=7, 6^4=9, 6^5=10,6^6=5,6^7=8,6^8=4,6^9=2\] and \[6\times 2 = 1\]

OpenStudy (anonymous):

oh it got cut off! last one was \[6^9=2\] and \[2\times 6 \equiv 1(\text{mod}11)\]

OpenStudy (anonymous):

on the other had you probably could have come up with the "2" right away.

OpenStudy (anonymous):

right ! thanks very much!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

I found that Euclidean algorithm can work it out as well, say \[6b \equiv 1(\mod 11)\] we have 5 = 11 - 6 * 1 1 = 6 - 5 * 1 and then 1 = 6 - 5 * 1 = 6 - (11 - 6 * 1) * 1 = 6 * 2 - 11 and we obtain \[b = 2\] in the case of \[130b \equiv 1 (\mod 11)\] we have 9 = 130 - 11 * 11 2 = 11 - 9 * 1 1 = 9 - 2 * 4 and then 1 = 9 - 2 * 4 = 9 - (11 - 9) * 4 = 9 * 5 - 11 * 4 = (130 - 11 * 11) * 5 - 11 * 4 = 130 * 5 - 11 * 59 and we obtain b = 5 besides, \[130b \equiv 9b \equiv 1(\mod 11) \] may make it easier :)

OpenStudy (anonymous):

ps: how to type (mod 11) instead of \[(\mod 11)\] isn't it "(\mod 11)" in slash blankets?